The distance between five consecutive crest on a transverse wave is 12m.what is the wavelength of the wave?
1
Expert's answer
2019-06-11T10:11:33-0400

the wavelength of the wave "\lambda=\frac{l}{n}"

"\lambda = \frac{12*2}{5} = 4.8 m"